OpenStudy (anonymous):

..HELP!!! D:...2 questions. 1]How many solutions are there for the equation? x2 + 16 = 8x anwsers: :] 2 :] 1 :] 0 2]How many solutions are there for the equation? x2 + 9x = -7 anwsers: same as the first question.

OpenStudy (anonymous):

There will always be two solutions.

OpenStudy (anonymous):

for both?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

any quadratic equation have two and only two solutions

OpenStudy (anonymous):

Oh Ok. Thank you.

OpenStudy (campbell_st):

1. there is 1 solution but its called a repeated solution x^2 - 8x + 16 = (x-4)^2 2. rewrite the quadratic x^2 + 9x - 7 = 0 this will have two unequal solutions but it needs the general quadratic formula to find the solutions